OpenStudy (anonymous):

On a particular map, every 1 16 inch represents 1 mile. How many miles does 312 inches represent?

OpenStudy (tkhunny):

1/16 " ==> 1 mile 1/8 " ==> 2 miles 1/4 " ==> 4 miles 1/2 " ==> 8 miles 1 " ==> 16 miles Are we getting there? Can you think of a faster way?

OpenStudy (anonymous):

I think I have to use like the proportional formula or something .

OpenStudy (tkhunny):

Good answer: \(312\;in\cdot \dfrac{1\;mi}{\dfrac{1}{16}\;in} = \)

OpenStudy (anonymous):

It says the answer is supposed to be 56 mi

OpenStudy (tkhunny):

56 mi * (1/16 in / 1 mi) = 3.5 in Are you sure we're not missing some decimals around here, smoewhere?